OpenStudy (anonymous):

5-(9-7)=(5-9)-7 is an example of the associative property. True or False ?

OpenStudy (anonymous):

FALSE

ganeshie8 (ganeshie8):

True. given statement is mathematically incorrect. but its an example of associative property anyway. you are just shifting the brackets here.. tricky one :)

OpenStudy (anonymous):

WHAT DO YOU MEAN?@ganeshie8

ganeshie8 (ganeshie8):

below is an example of mathematically correct associative property : a+ (b+ c) = (a + b) + c below is an example of mathematically incorrect associative property : a - (b - c) = (a - b) - c hmm still im not completely sure. but 99% i think its associative... not sure though !!

OpenStudy (anonymous):

'+'and'-' are just symbols,right?